It takes 5 seconds for sound to travel 1 mile through air. It takes only 1 second for sound to travel 1 mile through water. How many minutes faster will a sound 90 miles away travel through water than air?

Answer by Alan3354(69443)

It takes 5 seconds for sound to travel 1 mile through air. It takes only 1 second for sound to travel 1 mile through water. How many minutes faster will a sound 90 miles away travel through water than air?
---------------
"minutes faster" makes no sense, minutes is not a unit of speed.
----
To find how many minutes sooner it would arrive, calculate the 2 times.
d = r*t
t = d/r
------
In water, r = 1 mi/sec
In air, r = 1/5 mi/sec
90 miles in air at 0.2 mi/sec --> 90/0.2 = 450 seconds